I am trying to figure out how to figure out the following... I am not sure how to change this into a quadratic formula to solve...
x^2=2/3x
Please help!!!

Answer by htmentor(1343)

I am trying to figure out how to figure out the following... I am not sure how to change this into a quadratic formula to solve...
x^2=2/3x
========================================
A quadratic formula is not needed since we can first factor the equation by pulling out an x:
x^2 = (2/3)x -> x^2 - (2/3)x = 0 -> x(x - 2/3) = 0
Since if either factor = 0 the whole expression = 0, the equation has solutions x = 0, and x = 2/3